OpenStudy (jennychan12):

Evaluate the integral (x(x^2-1)^1/2) from 0 to 1

OpenStudy (jennychan12):

\[\int\limits_{0}^{1} x(\sqrt{x^2+1}dx\] sorry minor mistake above

OpenStudy (jennychan12):

please no trig sub because i haven't learned it yet

OpenStudy (kainui):

Don't worry, you can just use substitution here like normal. Do you see anything in there that could possibly be its derivative in there? Don't forget the derivative of a constant is 0.

OpenStudy (jennychan12):

ohh i see it now. u = x^2+1 du = 2x then plug all that stuff in.

OpenStudy (kainui):

Yup. =D